# TuneNSearch: a hybrid transfer learning and local search approach for solving vehicle routing problems

### v2

@@ Abstract
This paper introduces TuneNSearch, a hybrid transfer learning and local search approach for addressing different variants of the vehicle routing problem (VRP). Recently, multi-task learning has gained much attention for solving VRP variants. However, this adaptability often compromises the performance of the models. To address this challenge, we first pre-train a reinforcement learning model on the multi-depot VRP (MDVRP), followed by a fine-tuning phase to adapt it to different variants. In more detail, TuneNSearch employs, in the first stage, a Transformer-based architecture, augmented with a residual edge-graph attention network to capture the impact of edge distances and residual connections between layers. This architecture allows for a more precise capture of graph-structured data, improving the encoding of VRP’s features. After inference, our model is also coupled with a second stage composed of a local search algorithm, which yields substantial performance gains with small computational overhead added. Results show that TuneNSearch outperforms many existing state-of-the-art models trained for each VRP variant. The inherent complexity of the MDVRP enables the pre-trained model to generalize effectively to single-depot variants, achieving performance comparable to models trained specifically on single-depot instances. Moreover, it also demonstrates strong performance on multi-depot variants — a capability that models pre-trained exclusively on single-depot problems do not possess. Finally, through numerous experiments on randomly generated instances and benchmark public datasets, we show that our approach demonstrates strong generalization, achieving high performance across different tasks, distributions and problem sizes, thus addressing a long-standing gap in the literature.
@@
Vehicle Routing Problem , Combinatorial Optimization , Learning to Route , Local Search
journal: Elsevier
@@ 1 Introduction
The combinatorial characteristics of the VRP and its variants make these problems very difficult to solve optimally. Their inherent NP-hardness and intractability often render exact methods impractical, especially when solving large-scale problems with complicated constraints. Alternatively, meta-heuristics rely on search techniques to explore the solution space more efficiently. In this field, numerous methods have achieved notable performance on different routing problems, including LKH-3 (Helsgaun, 2017), hybrid genetic search (Vidal, 2022) and PyVRP (Wouda et al., 2024). However, despite presenting good results when solving complicated problems, compared with exact approaches, even the most efficient meta-heuristics still face significant computational burden as problem size grows. As a result, their applicability is limited in real-world cases, where rapid and precise decision-making is critical (Li et al., 2023).
@@
While neural-based methods demonstrate promising results, they are often tailored for specific types of routing problems, similar to meta-heuristics, which are typically designed for particular cases. This specialization limits their ability to generalize across different problem variants. Recent research efforts have sought to address this challenge. For instance, Liu et al. (2024) proposed a multi-task learning method that utilizes attribute composition to solve different VRP variants. Meanwhile, Zhou et al. (2024a), explored a different approach, using a mixture-of-experts model to enhance cross-problem generalization. Although these studies offer valuable insights, they still fall short in matching the performance of neural-based models that are specifically trained for each VRP variant. Their improved adaptability is noteworthy, however, it comes at the expense of solution quality, making them less effective than specialized models. In contrast, Lin et al. (2024) explored a fine-tuning approach that adapts a backbone model, pre-trained on a standard TSP, to effectively solve other variants. While this method outperforms models trained independently for each variant, it requires a similar computational effort and a comparable number of training epochs in the fine-tuning phase, making it overall less computationally efficient.
@@
Motivated by this challenge, we introduce TuneNSearch, a novel fine-tuning transfer learning approach designed to tackle various routing problems. Transfer learning is a machine learning technique that allows a model trained on one task to be adapted for a different but related task (Pan and Yang, 2010). In this way, previously acquired knowledge can be leveraged to improve the learning performance and reduce training time. Specifically, we propose pre-training a model on the MDVRP, a variant that has been largely overlooked by existing neural-based methods. Through exploiting the complexity inherent to the MDVRP, our goal is for the pre-trained model to not only generalize effectively to single-depot variants, but also multi-depot scenarios — a capability that models pre-trained exclusively on single-depot problems lack. Given that existing neural-based methods often struggle with generalization, particularly for larger instances, our approach features a hybridization combining machine learning with a high-performance local search algorithm. Therefore, a two-stage process is proposed, where an efficient local search refines solutions after the initial inference stage. As a result, TuneNSearch achieves significant improvements in generalization, substantially reducing the performance gaps compared to existing methods, such as Kwon et al. (2020), Li et al. (2024) and Zhou et al. (2024a). The key contributions of this work include:
@@
    Development of a novel transfer learning method that is based on pre-training on the MDVRP, designed to adapt effectively to various VRP variants. Unlike simpler variants, the MDVRP incorporates a multi-depot structure in which vehicles can depart from multiple locations, thereby allowing the model to capture richer and more complex representations. Computational results show that this approach facilitates a more efficient generalization to solve both single- and multi-depot VRP variants.
@@
    Proposal of a novel learning framework for solving the MDVRP. Our method employs a Transformer architecture (Vaswani et al., 2017) using the policy optimization with multiple optima (POMO) approach from Kwon et al. (2020). The proposed framework also integrates a residual edge-graph attention network (E-GAT), similar to Lei et al. (2022), to capture the impact of edge information and residual connections between consecutive layers. Unlike Lei et al. (2022), who applied the residual E-GAT encoder to the attention model from Kool et al. (2019), we extend it to POMO and the MDVRP. This extension enables a more accurate calculation of attention coefficients by incorporating edge distances during the feature encoding process. Our model outperforms the multi-depot multi-type attention (MD-MTA) approach proposed by Li et al. (2024). Specifically, in instances with 50 nodes, we reduced the performance gap by a factor of 6, and on instances with 100 nodes, by a factor of more than 2.
@@
    Improvement of the inference process by integrating an efficient local search method. This local search algorithm employs a set of different operators to improve the solutions obtained by the E-GAT model, resulting in significant performance gains with small additional computational cost. As a result, the proposed method achieves new state-of-the-art performance among neural-based approaches. On 100-node instances, it obtains an average gap of less than 4% across all VRP variants compared to PyVRP (Wouda et al., 2024), while requiring only a fraction of the computational time.
@@
    To verify the effectiveness of TuneNSearch, we solve numerous large-scale instances from CVRPLIB and TSPLIB. TuneNSearch demonstrates robust cross-distribution, cross-size and cross-task generalization, consistently exceeding other neural-based state-of-the-art methods in nearly all tested instances.
@@ 2.1 Solving VRPs with exact methods and meta-heuristics
Over the last few decades, a variety of methods have been proposed to address routing problems, including exact methods and meta-heuristic algorithms. Exact methods are fundamentally limited in their ability to guarantee optimal solutions within polynomial time. As a result, their application is typically limited to small- and medium-sized problem instances. The development and application of exact methods for routing problems has been reviewed by Baldacci et al. (2012) and Zhang et al. (2022). Most of these methods involve techniques such as branch-and-cut, dynamic programming and set partitioning formulations. More recently, Pessoa et al. (2020) introduced a generic solver based on a branch-cut-and-price algorithm capable of solving different routing problem variants. Nevertheless, the method remains computationally intractable for instances involving more than a few hundred nodes.
@@
Alternatively, meta-heuristic algorithms are far more prevalent than exact methods in the literature (Braekers et al., 2016). Unlike exact approaches, meta-heuristics do not guarantee optimal solutions, as a complete search of the solution space cannot be proven. However, these methods tend to be more efficient, utilizing advanced exploration techniques to find high-quality solutions — often reaching optimal solutions — in significantly less time. The first meta-heuristic method for the MDVRP was proposed by Renaud et al. (1996), introducing a tabu search algorithm that constructs an initial solution by assigning each customer to its nearest depot. The approach consists of three phases: fast-improvement, intensification and diversification. Marinakis and Marinaki (2010) combined a genetic algorithm with particle swarm optimization for the CVRP, allowing individual solutions within the population to evolve throughout their lifetime. This approach reduced computational time and improved scalability to larger problem instances compared to other meta-heuristics. Helsgaun (2017) introduced LKH-3, a heuristic method capable of solving various routing problem variants. It transforms the problems into a constrained TSP and utilizes the LKH local search (Helsgaun, 2000) to effectively explore the solution space. Silva et al. (2019) designed a multi-agent meta-heuristic framework combined with reinforcement learning for solving the VRP with time windows. In this framework, each meta-heuristic is represented as an autonomous agent that collaborates with others, enhancing solution quality through cooperative behavior. Vidal (2022) presented a hybrid genetic search for the CVRP and introduced the new Swap* operator. Rather than swapping two customers directly in place, this operator proposes exchanging two customers from different routes by inserting them into any position on the opposite route. This work has become a standard in CVRP literature, achieving state-of-the-art performance across various benchmarks. Kalatzantonakis et al. (2023) presented a hybrid approach between reinforcement learning and a variable neighborhood search for the CVRP, utilizing different upper confidence bound algorithms for adaptive neighborhood selection. More recently, Wouda et al. (2024) introduced PyVRP, an open-source VRP solver package. PyVRP offers a high-performance implementation of the hybrid genetic search algorithm (Vidal, 2022) and supports extensive customization options, making it suitable for a variety of VRP variants. Lastly, OR-Tools (Furnon and Perron, 2024) is a general-purpose optimization toolkit designed to solve a wide range of combinatorial problems. It includes a robust routing library capable of addressing various VRP variants, offering greater versatility than the previously cited works. OR-Tools uses a first-solution heuristic to generate an initial solution, followed by a guided local search to iteratively improve it. Subsequently, it employs a constraint programming approach to check whether the solution satisfies all specified constraints.
@@ 2.2 Neural-based combinatorial optimization for VRPs
Although meta-heuristics are generally more computationally efficient than exact methods, their execution time still increases significantly with the instance size. In this sense, neural-based methods surged in recent years as an alternative to solve combinatorial problems (Bengio et al., 2021; Mazyavkina et al., 2021). By recognizing patterns in data, these methods can learn policies, obtaining high-quality solutions in polynomial time, even for large and hard to solve instances.
Currently, there exists two main categories of neural methods for solving routing problems: construction-based and improvement-based methods. Construction-based methods learn a policy to construct solutions incrementally, starting from an empty set and building routes step by step in an autoregressive manner. Vinyals et al. (2015) introduced Pointer Networks, a sequence-to-sequence model that addressed the problem of variable-sized outputs by using a ‘pointer’ mechanism to select elements from the input sequence as the output. Their approach was applied to solve the TSP and trained using supervised learning, being an early demonstration of the potential of neural networks for combinatorial optimization. Later, Bello et al. (2017) built on this approach by using a similar model architecture but training it with reinforcement learning. This eliminated the need for (near)-optimal labels and led to improved performance over Pointer Networks. Nazari et al. (2018) extended the architecture of Pointer Networks to handle dynamic elements in problems. Their approach proved effective in solving more challenging combinatorial problems, such as the stochastic VRP and VRP with split deliveries. Kool et al. (2019) made a significant contribution to recent literature by proposing an attention model that utilizes a Transformer architecture (Vaswani et al., 2017). Trained using reinforcement learning, their method outperformed previous methods across a variety of combinatorial problems, representing a major advancement in the field. Kwon et al. (2020) introduced POMO, a reinforcement learning approach which leverages solution symmetries to improve results when compared to the attention model. POMO also introduced an instance augmentation technique, which reformulates a given problem by applying transformations — such as flipping or rotating the Euclidean map of node coordinates — to generate alternative instances that lead to the same solution. This technique forces the exploration of a wider range of potential solutions, enhancing model performance during inference. These works can be considered the backbone of the published research on routing problems, providing inspiration for a wide variety of subsequent studies (Bi et al., 2025; Chalumeau et al., 2023; Fitzpatrick et al., 2024; Grinsztajn et al., 2023; Kim et al., 2022; Kwon et al., 2021; Lei et al., 2022; Lin et al., 2024; Luo et al., 2023; Pirnay and Grimm, 2024; Xin et al., 2020; Zhou et al., 2023, 2024b).

### v3

@@ Abstract
This paper introduces TuneNSearch, a hybrid transfer learning and local search approach for addressing diverse variants of the vehicle routing problem (VRP). Our method uses reinforcement learning to generate high-quality solutions, which are subsequently refined by an efficient local search procedure. To ensure broad adaptability across VRP variants, TuneNSearch begins with a pre-training phase on the multi-depot VRP (MDVRP), followed by a fine-tuning phase to adapt it to other problem formulations. The learning phase utilizes a Transformer-based architecture enhanced with edge-aware attention, which integrates edge distances directly into the attention mechanism to better capture spatial relationships inherent to routing problems. We show that the pre-trained model generalizes effectively to single-depot variants, achieving performance comparable to models trained specifically on single-depot instances. Simultaneously, it maintains strong performance on multi-depot variants, an ability that models pre-trained solely on single-depot problems lack. For example, on 100-node instances of multi-depot variants, TuneNSearch outperforms a model pre-trained on the CVRP by 44%. In contrast, on 100-node instances of single-depot variants, TuneNSearch performs similar to the CVRP model. To validate the effectiveness of our method, we conduct extensive computational experiments on public benchmark and randomly generated instances. Across multiple CVRPLIB and TSPLIB datasets, TuneNSearch consistently achieves performance deviations of less than 3% from the best-known solutions in literature, compared to 6–25% for other neural-based models, depending on problem complexity. Overall, our approach demonstrates strong generalization to different problem sizes, instance distributions, and VRP formulations, while maintaining polynomial runtime complexity despite the integration of the local search algorithm.
@@
Vehicle Routing Problem , Combinatorial Optimization , Reinforcement Learning , Local Search
journal: Computers & Operations Research
@@ 1 Introduction
The combinatorial characteristics of the VRP and its variants make these problems very difficult to solve optimally. Their inherent NP-hardness and intractability often render exact methods impractical, especially when solving large-scale problems with complicated constraints. Alternatively, meta-heuristics rely on search techniques to explore the solution space more efficiently. Historically, meta-heuristics can be classified into single-solution based and population-based methods (Laporte, 2009). Single-solution based methods start from an individual solution and iteratively improve it through local search processes. They are relatively simple to implement and exhibit strong intensification, since they focus on promising areas of the solution search space. Examples of single-solution based methods include tabu search (Gendreau et al., 1994), simulated annealing (Osman, 1993), variable neighborhood search (Mladenović and Hansen, 1997) and adaptive large neighborhood search (Røpke and Pisinger, 2006). Their main limitation, however, is that they have a high risk of getting trapped in local minimum regions. In this regard, population-based algorithms show better exploration and diversification, since they maintain multiple solutions simultaneously, that is, a population of solutions. Common population-based methods include genetic algorithms (Baker and Ayechew, 2003), ant colony optimization (Dorigo and Gambardella, 1997) and particle swarm optimization (Ai and Kachitvichyanukul, 2009).
To leverage the strengths of both approaches, many state-of-the-art VRP methods combine single-solution and population-based strategies. One of the most prominent examples is the hybrid genetic search (Vidal, 2022), which integrates a local search method at each iteration of a genetic algorithm to improve the quality of offspring solutions. It has achieved state-of-the-art performance, outperforming multiple other well-known methods. More recently, Wouda et al. (2024) proposed PyVRP, a computationally efficient open-source implementation of the hybrid genetic search, designed to handle a range of VRP variants, instead of being tied to a single problem type. Despite presenting good results when solving complicated problems, compared with exact approaches, even the most efficient meta-heuristics still face significant computational burden as problem size grows. As a result, their applicability is limited in real-world cases, where rapid and precise decision-making is critical (Li et al., 2023).
@@
While neural-based methods demonstrate promising results, they are often tailored for specific types of routing problems, similar to meta-heuristics, which are typically designed for particular cases. This specialization limits their ability to generalize across different problem variants. Recent research efforts have sought to address this challenge. For instance, Liu et al. (2024) proposed a multi-task learning method that utilizes attribute composition to solve different VRP variants. Meanwhile, Zhou et al. (2024a), explored a different approach, using a mixture-of-experts model to enhance cross-task generalization. Although these methods improve flexibility, they typically require extensive training (up to 100 million instances, or 5+ days of training time, depending on the hardware). Moreover, their performance often falls short when compared to models trained specifically for individual variants, especially on more complex tasks. In contrast, Lin et al. (2024) explored a fine-tuning approach that adapts a backbone model, pre-trained on a standard TSP, to effectively solve other variants. While this method outperforms models trained independently for each variant, the fine-tuning phase demands a computational effort and training duration comparable to training a new model from scratch. Furthermore, existing multi-task and transfer learning approaches tend to focus exclusively on single-depot variants, overlooking the MDVRP, which more accurately reflects real-world logistics applications.
@@
Motivated by this challenge, we introduce TuneNSearch, a hybrid framework that combines transfer learning with an efficient local search procedure to solve different VRP variants. Transfer learning is a machine learning technique that allows a model trained on one task to be adapted for a different but related task (Pan and Yang, 2010). In this way, previously acquired knowledge can be leveraged to improve the learning performance and reduce training time. Specifically, we propose pre-training our model on the MDVRP, a rich and challenging problem setting that has been underexplored in neural approaches. By exploiting the complexity inherent to the MDVRP, our goal is for the model to adapt well to both single-depot and multi-depot scenarios, a capability missing in models pre-trained exclusively on single-depot data. Given that existing neural-based methods often struggle with generalization, particularly for larger instances, our approach features a hybridization combining machine learning with a high-performance local search algorithm. Therefore, a two-stage process is proposed, where an efficient local search refines solutions after the initial inference stage. As a result, TuneNSearch achieves significant improvements in generalization, substantially reducing the performance difference compared to existing methods, such as Kwon et al. (2020), Li et al. (2024) and Zhou et al. (2024a). The key contributions of this work include:
@@
    Improvement of the inference process by integrating an efficient local search method. This local search algorithm employs a set of different operators to improve the solutions obtained by the reinforcement learning model, resulting in significant performance gains with small additional computational cost. As a result, the proposed method achieves new state-of-the-art performance among neural-based approaches.
@@
    Development of a transfer learning method that begins with pre-training on the MDVRP, designed to adapt effectively to various VRP variants. Unlike simpler variants, the MDVRP incorporates a multi-depot structure in which vehicles can depart from multiple locations, thereby allowing the model to capture richer and more complex representations. Computational results show that this approach facilitates a more efficient generalization to solve both single- and multi-depot VRP variants.
@@
    Proposal of a novel learning framework to solve the MDVRP. Our method employs a Transformer architecture (Vaswani et al., 2017) using the policy optimization with multiple optima (POMO) (Kwon et al., 2020). To better capture the complex spatial and relational structure of MDVRP instances, we integrate a residual edge-aware graph attention network (E-GAT), inspired by Lei et al. (2022), but uniquely adapted to work within the POMO framework and extended to the multi-depot setting. This allows the model to incorporate edge distance information directly into the attention mechanism, leading to more accurate representations. Our method achieves substantial empirical improvements, outperforming the multi-depot multi-type attention (MD-MTA) approach proposed by Li et al. (2024). Specifically, in 50-node instances, we reduced the relative deviation from the best-performing method in our experiments by a factor of 6, and on 100-node instances by a factor of more than 2.
@@
    To verify the effectiveness of TuneNSearch, we solve numerous large-scale datasets from CVRPLIB and TSPLIB. On average, the solutions for each dataset deviate by no more than 3% from the best-known solutions in literature. We also assessed the performance of our method on thousands of randomly generated instances. For 100-node instances, across all VRP variants, it obtains an average relative performance deviation of less than 4% compared to PyVRP’s hybrid genetic search (Wouda et al., 2024), while requiring only a fraction of the computational time. Overall, TuneNSearch demonstrates robust cross-distribution, cross-size and cross-task generalization, consistently exceeding other neural-based methods.
@@ 2.1 Solving VRPs with exact methods and meta-heuristics
Over the last few decades, a variety of methods have been proposed to address routing problems, including exact methods and meta-heuristic algorithms. Exact methods are fundamentally limited in their ability to guarantee optimal solutions within polynomial time. As a result, their application is typically limited to small- and medium-sized problem instances. The development and application of exact methods for routing problems has been reviewed by Baldacci et al. (2012) and Zhang et al. (2022). Most of these methods involve techniques such as branch-and-cut, dynamic programming and set partitioning formulations. More recently, Pessoa et al. (2020) introduced a generic solver based on a branch-cut-and-price algorithm capable of solving different routing problem variants. Nevertheless, the proposed method remains computationally intractable for instances involving more than a few hundred nodes.
Alternatively, meta-heuristic algorithms are far more prevalent than exact methods in the literature (Braekers et al., 2016). Unlike exact approaches, meta-heuristics do not guarantee optimal solutions, as a complete search of the solution space cannot be proven. However, these methods tend to be more efficient, utilizing advanced exploration techniques to find high-quality solutions, and often even reaching optimal ones in significantly less time. Gendreau et al. (1994) proposed TABUROUTE, one of the first effective tabu search methods for routing problems. Specifically, the authors addressed the CVRP under route distance limit constraints. During the search process, they allow temporary violations of constraints, managing them with dynamically adjusted penalties in the objective function, improving exploration compared to methods that stay strictly within feasible regions of the search space. Subsequently, Renaud et al. (1996) presented one of the most influential meta-heuristic methods for the MDVRP, designing a tabu search algorithm that constructs an initial solution by assigning each customer to its nearest depot. The approach consists of three phases: fast-improvement, intensification and diversification. The first phase aims at rapidly improving the incumbent solution using different search operators. The intensification phase then focuses on exploring high-potential areas of the search space, while the last one introduces controlled perturbations to escape local minimum and enable broader exploration. Mladenović and Hansen (1997) later introduced the variable neighborhood search meta-heuristic, a method that systematically changes neighborhood structures during the search process. Each time the algorithm explores a different neighborhood of the incumbent solution, it applies a local search to find a local optimum within that neighborhood. This local search serves to exploit the current region of the solution space, while the systematic change of neighborhoods allows the algorithm to escape local minimum. Although initially tested on the TSP, the method’s effectiveness have established it as a standard within combinatorial optimization literature. After, Baker and Ayechew (2003) proposed one of the first genetic algorithms for the classic VRP, offering a new perspective on solving routing problems. The authors also developed a hybrid method that incorporates neighborhood search mechanisms within the genetic algorithm to improve the balance between exploration and intensification. This hybridization of single-solution based and population-based methods would later become a widely established practice in VRP research. Computational results demonstrated the effectiveness of the genetic algorithm, showing performance competitive with other state-of-the-art methods at the time, including a tabu search and a simulated annealing algorithm. Ai and Kachitvichyanukul (2009) presented a particle swarm optimization algorithm for the VRP with simultaneous delivery and pickup. Tests on multiple benchmark datasets showed that their method performed competitively against other algorithms and even achieved new best-known solutions for various benchmark instances. Although not directly based on the particle swarm optimization method of Ai and Kachitvichyanukul (2009), Marinakis and Marinaki (2010) explored a related idea by hybridizing a genetic algorithm with particle swarm optimization for the CVRP, allowing individual solutions within the population to evolve. This approach reduced computational time and improved scalability to larger problem instances compared to other meta-heuristics.
@@
Helsgaun (2017) introduced LKH-3, a heuristic method capable of solving various routing problem variants. It transforms the problems into a constrained TSP and utilizes the LKH local search from Helsgaun (2000) to effectively explore the solution space. Silva et al. (2019) designed a multi-agent meta-heuristic framework combined with reinforcement learning for solving the VRP with time windows. In this framework, each meta-heuristic is represented as an autonomous agent that collaborates with others, enhancing solution quality through cooperative behavior. Vidal (2022) presented a hybrid genetic search for the CVRP and introduced the new Swap* operator. Rather than swapping two customers directly in place, this operator proposes exchanging two customers from different routes by inserting them into any position on the opposite route. By combining elements from both single-solution based and population-based meta-heuristics, this work has become a standard in CVRP literature, achieving state-of-the-art performance across various benchmarks. Kalatzantonakis et al. (2023) presented a hybrid approach between reinforcement learning and a variable neighborhood search for the CVRP, utilizing different upper confidence bound algorithms for adaptive neighborhood selection. More recently, Wouda et al. (2024) introduced PyVRP, an open-source VRP solver package. PyVRP offers a high-performance implementation of the hybrid genetic search algorithm (Vidal, 2022) and supports extensive customization options, making it suitable for a variety of VRP variants. Lastly, OR-Tools (Furnon and Perron, 2024) is a general-purpose optimization toolkit designed to solve a wide range of combinatorial problems. It includes a robust routing library capable of addressing various VRP variants, offering greater versatility than the previously cited works. OR-Tools uses a first-solution heuristic to generate an initial solution, followed by a guided local search to iteratively improve it. Subsequently, it employs a constraint programming approach to check whether the solution satisfies all specified constraints.
@@ 2.2 Neural-based combinatorial optimization for VRPs
Neural-based methods surged in recent years as an alternative to solve combinatorial problems (Bengio et al., 2021; Mazyavkina et al., 2021). By recognizing patterns in data, these methods can learn policies, obtaining high-quality solutions in polynomial time, even for large and hard to solve instances.
As previously discussed, neural-based approaches can be categorized into construction or improvement methods. In what follows, we focus first on construction-based methods, which learn policies to incrementally build a solution. Vinyals et al. (2015) introduced Pointer Networks, a sequence-to-sequence model that addressed the problem of variable-sized outputs by using a ‘pointer’ mechanism to select elements from the input sequence as the output. Their approach was applied to solve the TSP and trained using supervised learning, being an early demonstration of the potential of neural networks for combinatorial optimization. Later, Bello et al. (2017) built on this approach by using a similar model architecture but training it with reinforcement learning. This eliminated the need for (near)-optimal labels and led to improved performance over Pointer Networks. Nazari et al. (2018) extended the architecture of Pointer Networks to handle dynamic elements in problems. Their approach proved effective in solving more challenging combinatorial problems, such as the stochastic VRP and VRP with split deliveries. Kool et al. (2019) made a significant contribution to recent literature by proposing an attention model that utilizes a Transformer architecture (Vaswani et al., 2017). Trained using reinforcement learning, their method outperformed previous methods across a variety of combinatorial problems, representing a major advancement in the field. Kwon et al. (2020) introduced POMO, a reinforcement learning approach which draws on solution symmetries to improve results when compared to the attention model. POMO also introduced an instance augmentation technique which reformulates a given problem by applying transformations, such as flipping or rotating the Euclidean map of node coordinates, to generate alternative instances that lead to the same solution. This technique forces the exploration of a wider range of potential solutions, enhancing model performance during inference. These works can be considered the backbone of the published research on routing problems, providing inspiration for a wide variety of subsequent studies (Bi et al., 2025; Chalumeau et al., 2023; Fitzpatrick et al., 2024; Grinsztajn et al., 2023; Kim et al., 2022; Kwon et al., 2021; Lei et al., 2022; Lin et al., 2024; Luo et al., 2023; Pirnay and Grimm, 2024; Xin et al., 2020; Zhou et al., 2023, 2024b).